# APRIL: Active Partial Rollouts in Reinforcement Learning to tame long-tail generation

### v1

@@ 3.2.3 Analysis of Percentage of Partial Rollouts and Rollout Length
How much off-policy data does APRIL introduce?
A key feature of our partial rollout method is that tokens from incomplete trajectories are cached and reused in subsequent steps, as illustrated in Figure 2. This process inherently introduces off-policy data. We analyzed the rollout composition of our training batches and found this effect to be substantial. As shown in Figure 6, in a typical step with over-sampling, nearly all unfinished responses among the $256$ additional requests are partially generated and then resumed in the next iteration. Tokens carried over from these truncated instances constitute approximately 40% of the tokens in the subsequent steps. Despite this high proportion, APRIL remains remarkably stable throughout training and even achieves slightly higher final accuracy. This observation suggests that incorporating rollouts from multiple policy models may diversify the data and enhance the robustness of training.
In the previous APRIL framework, we discussed partial rollouts at the batch level, where rollout lengths exhibit substantial variability across different input instances within a batch. To improve stability and facilitate finer-grained control of rollouts for each instance, we adopt an instance-level approach. Specifically, in our implementation, each input instance forms a group comprising all rollout responses associated with that instance (e.g., n_samples_per_prompt=8=8). The completed rollouts are retained in the data buffer until the remaining rollouts are finished, after which they are collectively used to compute the reward for updating the policy model.
A potential concern is that instance-level granularity might introduce an intra-group long-tail effect, where generated rollouts for a single instance vary significantly in length. To examine this, we compare rollout-length variability at two levels of granularity: batch level and instance level.
At the batch level, the standard deviation $\sigma_{batch-level}$ of response lengths across all prompts within an iteration is approximately 4,0004{,}000-4,5004{,}500 tokens across datasets (Figure 7). In contrast, at the instance level (within the same input instance), the $\sigma_{instance-level}$ is typically less than 1,5001{,}500 tokens for GRPO algorithm, and 2,4002{,}400 tokens for DAPO algorithm(Figure 8). This pronounced gap indicates that responses sampled from the same input instance are considerably more homogeneous in length compared to responses sampled across different instances. Consequently, instance-level completion does not reintroduce a severe intra-group long-tail effect.
3.3 Hardware-Agnostic
To demonstrate the compatibility of APRIL and the general applicability, we merged APRIL into the most used open-source RL framework - slime and replicated our main experiment across multiple hardware platforms. The results in Figure 9 show consistent and comparable percentage improvements on NVIDIA H100 and AMD MI300. To enable execution on the MI300, we modified the memory management mechanism 888torch_memory_saver: https://github.com/fzyzcjy/torch_memory_saver/pull/43. We also have a HIP ROCm version that supports running on AMD GPUs already. to support a synchronous collated-model RL training framework, which is also the type of framework that we use. This demonstrates that APRIL is a generic system-level optimization that effectively addresses a fundamental bottleneck in the RL workflow. Since we have shown the results on AMD GPUs, we only show the results on NVIDIA H100 here.
(a) Throughput
(b) Evaluation Score
Figure 9: Throughput and evaluation scores on NVIDIA H200.
4 Related Work
In related works, extend our discussion by providing additional background and analyzing the system-level and algorithmic bottlenecks in existing RL frameworks. Specifically, we conduct a literature survey on RL framework architectures, existing rollout optimization techniques, and rollout categories to further clarify the motivation for designing APRIL.
4.1 RL Framework Architectures: Synchronous vs. Asynchronous
@@ 4.2 Rollout Optimization
Beyond the fine-grained improvements at the inference level, other system-level optimizations aim to address broader inefficiencies in the overall RL workflow. As previously discussed in Section 4.1, fully asynchronous RL (Noukhovitch et al., 2025) represents the most typical and intuitive approach to optimize RL workflows at this level by completely decoupling computational components. Within the synchronous RL settings, recent works such as SortedRL (Zhang et al., 2025) and RLHFuse (Zhong et al., 2025b) have explored system-level optimizations to mitigate these inefficiencies via length-aware scheduling and pipeline fusion mechanisms. The term partial rollout was first mentioned in the technical report for Kimi k1.5 (Kimi et al., 2025), where it was described as a technique to improve training efficiency for long-context RL. Specifically, it was conceptualized as ”sampling new trajectories by reusing a large chunk of previous trajectories, avoiding the cost to re-generate the new trajectories from scratch”. This concept was presented within the broader context of long context scaling and improved policy optimization. However, to the best of our knowledge, the specific mechanism, implementation, and its impact on system throughput and model convergence have not been detailed publicly. Our work implements a specific form of this concept as a preemptive scheduling mechanism designed for the iterative loop of RL training.
4.3 Algorithm: On-policy vs. Off-policy
@@
5 Conclusion
In this work, we presented APRIL, a simple yet effective approach to alleviating long-tail inefficiency in reinforcement learning for large language models. By actively reusing partial rollouts and avoiding wasted computation, APRIL substantially improves throughput and accelerates convergence without sacrificing accuracy. Beyond its empirical benefits, APRIL highlights the importance of jointly addressing system bottlenecks and algorithmic design in order to scale RL training to ever-larger models and tasks. As rollout generation continues to dominate the cost of RL, we believe APRIL provides a practical step toward more efficient and sustainable training pipelines. Looking forward, we anticipate that the principles behind APRIL will inspire future work on adaptive rollout strategies, tighter integration with inference engines, and more broadly, the design of RL frameworks that fully exploit modern hardware.

### v3

@@ A.3.3 Analysis of Percentage of Partial Rollouts and Rollout Length
A.4 Motivation and More Background
In the main paper, due to page limitations, we provided only a brief discussion to demonstrate the motivation. Here, we present a more comprehensive discussion by offering additional background and analyzing both system-level and algorithmic bottlenecks in existing RL frameworks. Specifically, we conduct a literature survey on RL framework architectures, existing rollout optimization techniques, and rollout categories to further clarify the motivation behind the design of APRIL.
A.4.1 RL Framework Architectures: Synchronous vs. Asynchronous
@@ A.4.2 Rollout Optimization
Beyond the fine-grained improvements at the inference level, other system-level optimizations aim to address broader inefficiencies in the overall RL workflow. As previously discussed in Section, fully asynchronous RL (Noukhovitch et al., 2025) represents the most typical and intuitive approach to optimize RL workflows at this level by completely decoupling computational components. Within the synchronous RL settings, recent works such as SortedRL (Zhang et al., 2025) and RLHFuse (Zhong et al., 2025b) have explored system-level optimizations to mitigate these inefficiencies via length-aware scheduling and pipeline fusion mechanisms. The term partial rollout was first mentioned in the technical report for Kimi k1.5 (Kimi et al., 2025), where it was described as a technique to improve training efficiency for long-context RL. Specifically, it was conceptualized as ”sampling new trajectories by reusing a large chunk of previous trajectories, avoiding the cost to re-generate the new trajectories from scratch”. This concept was presented within the broader context of long context scaling and improved policy optimization. However, to the best of our knowledge, the specific mechanism, implementation, and its impact on system throughput and model convergence have not been detailed publicly. Our work implements a specific form of this concept as a preemptive scheduling mechanism designed for the iterative loop of RL training.
A.4.3 Algorithm: On-policy vs. Off-policy